a ball is dropped from rest at a height of s=71.36 meters above the ground. calculate its speed before it hits the ground
1
Expert's answer
2020-07-13T11:41:11-0400

The potential energy of the ball at the height of s=71.36 is the following:


"E_p = mgs"

According to the energy conservation law, it will be equal to the kinetic energy of the ball before it hits the ground, which is given by the following expression:


"E_k = \dfrac{mv^2}{2}"

Thus:

"E_k = E_p\\
mgs = \dfrac{mv^2}{2}"

Expressing the spped, obtain:


"v^2 = 2gs\\
v = \sqrt{2gs}"

Substituting the numerical values, get:


"v = \sqrt{2\cdot 9.81m/s^2 \cdot 71.36m} \approx 18.46m/s"

Answer. 18.46 m/s.